Help please !
3 complete)D This Test: 40 pts Find the probability that in 200 tosses of a fair die, we will obtain at most 30 fives. Use the normal distribution to decimal places O A. 04936 OB. 0.2954 ° C. 0.1871 O D. 0.3229 Click to select your answer
Answer #1

probability of getting 5 = 1/6

X = number of time fives occur

X follow N (np , np*q)

here n = 200

N(200*1/6 , 200*1/6*5/6) = N ( 33.333333 , 27.7777777778 )

=P(X <= 30)

= P(X <= 30.5)

= P(Z < ( 30.5 - 33.333333)/ sqrt(27.7777777778))

= P(Z < -0.537587 )

=    0.2954

option B) is correct